two angles are complementary. the measure of one angle is 21 more than twice the measure of the other angle. find the measure of the angles.
1
Expert's answer
2012-09-07T10:14:25-0400
Sum of complementary angles is 90
Let measure of the second angle is x. Then measure of the first is 2x+21
So
2x+21+x=90
3x=90-21
x=69/3=23 - second angle
first 23*2+21=67